OpenStudy (anonymous):

The cost of renting a van for one day includes a flat rental fe plus a charge for each mile the van is driven when it is rented. A van that is driven 107 mle costs $97.15. A van that is driven 127 mils costs $106.15. What is the flat rental fee?

OpenStudy (anonymous):

127 miles

OpenStudy (jim766):

Do you know how to find eq of a line if you are given 2 points?

OpenStudy (anonymous):

yea. y=mx+b

OpenStudy (jim766):

They gave you 2 points (107, 97.50) (127, 106.15)

OpenStudy (anonymous):

i need to find the slope of that?

OpenStudy (anonymous):

what is it that i need to do?

OpenStudy (jim766):

Yeah , but note the first point should be (107, 97.15) Not what I put Y2-y1/x2-x1= slope to start

OpenStudy (anonymous):

can u list step by step what i have to do?

OpenStudy (jim766):

106.15-97.15/127-107 9/20 =.45. That is the slope

OpenStudy (jim766):

Now need to find. Y = mx + b Pick one of the points, (107. 15, 97.15) 97.15= .45(107.15) + b Solve for b and that will be y intercept

OpenStudy (anonymous):

ok. then what?

OpenStudy (jim766):

Sorry I mistyped again, duh 97.15= .45(107) + b 97.15= 48.15 + b -48.15. -48.15 49= b That is the flat fee. $49

OpenStudy (anonymous):

oh ok. thanku :)